OpenStudy (anonymous):

An orchestra consists of 25 members. The youngest member is 32 years old. She leaves and is replaced by a new member who is 30 years old. 1A : 2.0 POINTS By how much does the replacement change the average age of the members of the orchestra? [Note on answer format: Calculate your answer in years. If you think the average is unchanged, enter 0. If you think it goes up by 10.5 years, enter 10.5. If you think it goes down by 10.5 years, enter -10.5.] unanswered

OpenStudy (anonymous):

\[ \mu=\frac{32+\sum\limits_{k=2}^Na}{N}\\ \frac{\sum\limits_{k=2}^Na}{N}=\mu-{32\over N}\\ \mu'={30\over N}+\frac{\sum\limits_{k=2}^Na}{N}={30\over N}+\left(\mu-{32\over N}\right)\\ \mu'-\mu={30-32\over 25}={-2\over25}\\ \boxed{\mu'-\mu=-0.08\rm{years}} \]

OpenStudy (anonymous):

Its mean, we have to enter -0.08, this is 1A what about 1B?

Directrix (directrix):

http://openstudy.com/study#/updates/51513785e4b0d02faf5be2e9 What is question 1B?